# Text-Queried Audio Source Separation via Hierarchical Modeling

### v1

@@ 3 Sets 4000 10 1000
Additionally, we assess zero-shot audio separation performance on MUSIC dataset [8], which comprises 536 video recordings of individuals playing musical instruments across 11 classes (e.g., accordion, acoustic guitar, cello). Following [16], we downloaded 46 video recordings from the test split, randomly selected two clips from different instrument classes and mixed them into 10-second segments, producing 500 audio pairs. The average SNR in mixing is within [-5dB, 5dB].
@@ IV-B2 Objective metrics
We evaluate our separation performance using log spectral distance (LSD), peak signal-to-noise ratio (PSNR), and Kullback-Leibler divergence (KL), following the AudioLDM_eval audio generation evaluation pipeline [^2] https://github.com/haoheliu/audioldm_eval. LSD quantifies the difference between spectrograms of the predicted and target samples. PSNR measures the logarithmic ratio of the maximum possible signal power to the mean squared error between the predicted and target signals. KL divergence measures the similarity between the predicted and target audio with the label calculated by an audio tagging model.
To further evaluate the semantic correction of predicted audios, we leverage the CLAP score [62], which calculates audio-text similarity based on CLAP models [26, 25]. The MSCLAP [26] is chosen for this measurement in our experiment. As CLAP tends to poorly capture temporal semantics and complex multi-source audio semantics [63], we further introduce another semantic score, AFSim, by calculating a cosine similarity between predicted and target signals on their semantic embeddings based on large audio language models (AudioLLM). Specifically, we leverage Audio Flamingo [64], an audio LLM with advanced audio understanding capabilities, extract the feature from the penultimate (last second) layer by using the captioning prompt, and perform a mean-pooling to get the final embedding $AF\in\mathbb{R}^{1\times 2048}$, which is taken as the semantic embedding for our AFSim score measurement. The superior semantic representation capability of the $AF$ over CLAP is demonstrated in Table VI in audio captioning evaluation.
@@ IV-C Implementation details
We train the local-semantic separation module with a 12-layer NAR transformer which has 8 attention heads, and a hidden dimenion of 768. During two-stage joint fine-tuning, we utilize the same training data and fintune the parameters of two NAR transformers, with other modules kept frozen. $P_{gt}$ is set to 0.1. The loss weights of $\mathcal{L}_{global}$ and $\mathcal{L}_{local}$ are set to 0.1 and 1.0, respectively.
The acoustic decoder includes two parts. The autoregressive transformer has 12 layers with 8 attention heads and a hidden dimension of 1024. We take TF-Codec with a bitrate of 6 kbps with a causal setting, similar to that in [50].
@@ V Results
In this section, we evaluate the performance of our HSM-TSS approach and its modules including global-local representations and neural audio codec.
@@ V-A Evaluation results on general audio
TABLE II: Objective evaluation of separation for general audio
Model Train.Data 3 Sets Clotho AudioCaps FSD50K
[…]
LASS-Net 17 2.577 3.170 16.42 2.713 1.696 18.21 2.446 2.276 17.64 2.474 4.528 15.17
CLIPSep 550 2.320 3.197 14.38 2.616 1.634 18.43 2.738 2.464 17.07 2.967 4.475 17.15
BiModalSep 670 1.634 3.118 19.49 1.819 1.613 19.71 1.441 2.256 19.44 1.789 4.428 18.89
AudioSep 14100 1.027 3.037 22.50 1.191 1.616 21.57 1.002 2.142 21.17 1.172 4.313 23.66
@@
Ours_single 600 0.999 2.878 25.50 1.144 1.395 23.80 0.960 1.916 23.49 0.948 4.197 27.11
Ours 600 0.924 2.848 25.77 1.089 1.378 24.05 0.910 1.884 23.93 0.889 4.156 27.34
@@
Ours_single 600 1.079 2.869 25.50 1.184 1.383 23.98 1.148 1.919 23.66 1.069 4.209 26.65
Ours 600 1.007 2.852 25.66 1.112 1.373 23.94 1.105 1.908 23.88 1.041 4.179 27.00
TABLE III: Semantic evaluation for general audio
Model 3 Sets Clotho AudioCaps FSD50K
[…]
GT - 0.481 - 0.437 - 0.575 - 0.466
LASS-Net 0.615 0.268 0.590 0.314 0.641 0.264 0.592 0.223
CLIPSep 0.514 0.229 0.487 0.257 0.519 0.214 0.525 0.203
BiModalSep 0.678 0.420 0.674 0.385 0.713 0.501 0.631 0.387
AudioSep 0.730 0.357 0.720 0.401 0.752 0.376 0.699 0.305
@@
Ours_single 0.744 0.420 0.734 0.379 0.763 0.500 0.721 0.428
Ours 0.752 0.436 0.737 0.393 0.767 0.511 0.729 0.437
@@
Ours_single 0.738 0.419 0.730 0.377 0.755 0.486 0.718 0.420
Ours 0.741 0.428 0.733 0.380 0.763 0.492 0.720 0.425
@@
We compare the separation performance of our HSM-TSS with several text-queried audio separation methods, LASSNet [15], AudioSep [4], CLIPSep [16] and BiModalSep [17], which are all frequency-domain mask-based approaches. LASS-Net employs a pre-trained BERT as the text query encoder and ResUNet as the separation model, while AudioSep further integrates a CLAP text encoder as the query network and trains on a much larger dataset, yielding substantial performance gains. CLIPSep adopts CLIP as the query encoder and a Sound-of-Pixels (SOP)-based separation model, trained on approximately 500 hours of noisy audio-visual data from the VGGSound [24] dataset using hybrid vision-text supervision. BiModalSep introduces a weakly-supervised approach and leverages bi-modal semantic similarity via CLAP to align single-source language prompts with audio predictions, achieving robust separation without curated single-source data. We evaluate their performance using the official open-sourced models on our benchmarks.
@@
To show the effectiveness of our hierarchical modeling, we also compare with a single-separation-stage variant of our HSM-TSS, termed “Ours_single” in Table II. It removes the first global-semantic separation stage and only leverages the local-semantic separation stage. The text encoder is FLAN-T5 without Q-Audio in this variant. We also provide results for not only “extraction” as used in four compared methods but also “removal” task types that produce the same target audios.
@@
As shown in Table II, the proposed HSM-TSS, noted as “Ours”, outperforms both baseline methods and its single-separation-stage variant across various metrics. When compared to four methods, LASSNet, CLIPSep, AudioSep, and BiModalSep, our approach consistently achieves higher scores in the three metrics, especially for the 3 Sets benchmark, highlighting its effectiveness in separating target audio events from complex mixtures. In contrast to the single-separation-stage setting, our dual-stage separation allows for progressive refinement of audio features from coarse to fine, leading to more precise separation of target sounds, as demonstrated in Table II. We can also see that our model with “removal” instruction achieves comparable scores with “extraction”, showing the bidirectional operation capability of our method.
@@
Table III shows the evaluation results on semantic consistency of the output audios with instructions. For “removal” setting, we use the same audio description as “extraction” to calculate the CLAP score. We can see that our method consistently outperforms other methods and the single-separation-stage variant in both AFSim and CLAP scores, showing its superior capability to follow instructions with diverse audio descriptions. For all these comparisons, we only use the simple “extract” and “removal” instruction templates as the compared methods are not designed for open vocabulary separation instruction inputs but only target audio captions. Our method has good arbitrary instruction following capability, as shown in our demo page [^3] https://hsm-tss.github.io.
@@ V-B Zero-shot performance on unseen datasets
TABLE IV: Evaluation on zero-shot MUSIC dataset
[…]
GT - - - 0.490 -
LASS-Net 4.180 2.066 13.23 0.126 0.503
CLIPSep 2.179 2.554 14.07 0.372 0.657
BiModalSep 2.607 2.250 14.35 0.374 0.671
AudioSep 0.535 1.624 22.01 0.342 0.804
@@
Ours_single 0.585 1.400 23.22 0.463 0.799
Ours 0.501 1.375 23.85 0.470 0.811
V-C Visualization
V-C1 t-SNE visualization
To show how the global-semantic separation performs, we visualize the extracted features from this stage with t-SNE [65]. In Figure 5, each color shows a sound event class and we present the ground-truth global audio semantic feature and the separated output with different markers. It can be observed that our global-semantic separation model effectively extracts target audio features with good clustering.
@@ V-C Visualization
V-C2 Visualization of attention maps
@@
V-C3 Visualization of separation outputs
@@
V-D Hierarchical representations
V-D1 Global semantic feature evaluation
To evaluate the cross-modal alignment and semantic representation capbility of our Q-Audio, we follow the settings of LAION-CLAP [25] on AudioCaps and Clotho datasets. Table V shows the text-to-audio retrieval results. Compared to the widely used LAION-CLAP [25] and MSCLAP [26], Q-Audio achieves better performance, underscoring its capability in audio-language modeling. Table VI shows the audio captioning evaluation, where a linear mapping layer and a pretrained GPT2 [66] are used as the downstream model to generate audio captions. Here we only train the downstream model and keep audio feature extraction frozen. Although our Q-Audio is not pretrained on the captioning task as CLAP models, it still achieves higher scores on the three metrics, showing its excellent semantic representation capability.
TABLE V: Text-to-audio retrieval performance

### v2

@@ 3 Sets 4000 10 1000
Additionally, we assess zero-shot audio separation performance on MUSIC dataset [zhao2018sound], which comprises 536 video recordings of individuals playing musical instruments across 11 classes (e.g., accordion, acoustic guitar, cello). Following [dong2022clipsep], we downloaded 46 video recordings from the test split, randomly selected two clips from different instrument classes and mixed them into 10-second segments, producing 500 audio pairs. The average SNR in mixing is within [-5dB, 5dB].
@@ IV-B2 Objective metrics
We evaluate our separation performance using log spectral distance (LSD), peak signal-to-noise ratio (PSNR), and Kullback-Leibler divergence (KL), following the AudioLDM_eval audio generation evaluation pipeline [^2] https://github.com/haoheliu/audioldm_eval. LSD quantifies the difference between spectrograms of the predicted and target samples. PSNR measures the logarithmic ratio of the maximum possible signal power to the mean squared error between the predicted and target signals. KL divergence measures the similarity between the predicted and target audios with the probability distribution calculated by an audio tagging model. While metrics like SI-SDR are common, they are highly sensitive to perceptually irrelevant artifacts, e.g. phase shifts or subtle but continuous hums, which can be introduced by generative decoders.
To further evaluate the semantic correction of predicted audios, we leverage the CLAP score [xiao2024reference], which calculates audio-text similarity based on CLAP models [elizalde2023clap, wu2023large]. The MSCLAP [elizalde2023clap] is chosen for this measurement in our experiment. As CLAP tends to poorly capture temporal semantics and complex multi-source audio semantics [wu2023audio], we further introduce another semantic score, AFSim, by calculating a cosine similarity between predicted and target signals on their semantic embeddings based on large audio language models (AudioLLM). Specifically, we leverage Audio Flamingo [kong2024audio], an audio LLM with advanced audio understanding capabilities, extract the feature from the penultimate (last second) layer by using the captioning prompt, and perform a mean-pooling to get the final embedding $AF\in\mathbb{R}^{1\times 2048}$, which is taken as the semantic embedding for our AFSim score measurement. The superior semantic representation capability of the AudioFlamingoAudio\ Flamingo over CLAP is demonstrated in Table VII in audio captioning evaluation.
@@ IV-C Implementation details
We train the local-semantic separation module with a 12-layer NAR transformer which has 8 attention heads, and a hidden dimension of 768. During two-stage joint fine-tuning, we utilize the same training data and fintune the parameters of two NAR transformers, with other modules kept frozen. $P_{gt}$ is set to 0.1. The loss weights of $\mathcal{L}_{global}$ and $\mathcal{L}_{local}$ are set to 0.1 and 1.0, respectively.
The semantic-to-acoustic and acoustic decoder includes two parts. The autoregressive transformer has 12 layers with 8 attention heads and a hidden dimension of 1024. We take TF-Codec with a bitrate of 6 kbps in a causal setting, similar to that in [jiang2024universal].
@@ V Results
In this section, we evaluate the performance of our HSM-TSS approach and its modules including global-local representations and neural audio codec. For all these comparisons of separation performance, we only use the simple “extract” and “removal” instruction templates as the compared methods are not designed for open vocabulary separation instruction inputs but only target audio captions.
@@ V-A Evaluation results on general audio
TABLE II: Objective and semantic evaluation of separation for general audio (3 Sets and Clotho). “Ours_local” denotes a single-separation-stage variant using only local-semantic separation with FLAN-T5 text encoder. “Ours_local*” denotes a single-separation-stage variant using only local-semantic separation with Q-Audio text encoder. No global-semantic separation is used in these two variants.
Model Train.Data 3 Sets Clotho
(hrs) KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)} KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)}
GT - - - - - 0.481 - - - - 0.437
LASS-Net 17 2.577 3.170 16.42 0.615 0.268 2.713 1.696 18.21 0.590 0.314
CLIPSep 550 2.320 3.197 14.38 0.514 0.229 2.616 1.634 18.43 0.487 0.257
BiModalSep 670 1.634 3.118 19.49 0.678 0.420 1.819 1.613 19.71 0.674 0.385
AudioSep 14100 1.027 3.037 22.50 0.730 0.357 1.191 1.616 21.57 0.720 0.401
OmniSep 550 2.081 3.167 17.35 0.627 0.360 2.445 1.644 18.59 0.601 0.364
FlowSep 1680 1.580 3.100 19.19 0.711 0.428 1.366 1.654 20.28 0.718 0.378
@@
Ours_local 600 0.999 2.878 25.50 0.744 0.420 1.144 1.395 23.80 0.734 0.379
Ours_local* 600 0.997 2.879 25.51 0.741 0.421 1.135 1.383 23.95 0.730 0.381
Ours 600 0.924 2.848 25.77 0.752 0.436 1.089 1.378 24.05 0.737 0.393
@@
Ours_local 600 1.079 2.869 25.50 0.738 0.419 1.184 1.383 23.98 0.730 0.377
Ours_local* 600 1.061 2.866 25.53 0.734 0.420 1.182 1.381 23.99 0.728 0.376
Ours 600 1.007 2.852 25.66 0.741 0.428 1.112 1.373 23.94 0.733 0.380
TABLE III: Objective and semantic evaluation of separation for general audio (AudioCaps and FSD50K). “Ours_local” denotes a single-separation-stage variant using only local-semantic separation with FLAN-T5 text encoder. “Ours_local*” denotes a single-separation-stage variant using only local-semantic separation with Q-Audio text encoder. No global-semantic separation is used in these two variants.
Model Train.Data AudioCaps FSD50K
(hrs) KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)} KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)}
GT - - - - - 0.575 - - - - 0.466
LASS-Net 17 2.446 2.276 17.64 0.641 0.264 2.474 4.528 15.17 0.592 0.223
CLIPSep 550 2.738 2.464 17.07 0.519 0.214 2.967 4.475 17.15 0.525 0.203
BiModalSep 670 1.441 2.256 19.44 0.713 0.501 1.789 4.428 18.89 0.631 0.387
AudioSep 14100 1.002 2.142 21.17 0.752 0.376 1.172 4.313 23.66 0.699 0.305
OmniSep 550 2.174 2.401 18.03 0.640 0.403 1.897 4.398 19.25 0.627 0.326
FlowSep 1680 1.238 2.012 20.37 0.751 0.508 1.683 4.335 19.27 0.682 0.425
@@
Ours_local 600 0.960 1.916 23.49 0.763 0.500 0.948 4.197 27.11 0.721 0.428
Ours_local* 600 1.034 1.895 23.71 0.742 0.498 1.068 4.209 26.65 0.720 0.430
Ours 600 0.910 1.884 23.93 0.767 0.511 0.889 4.156 27.34 0.729 0.437
@@
Ours_local 600 1.148 1.919 23.66 0.755 0.486 1.069 4.209 26.65 0.718 0.420
Ours_local* 600 1.152 1.917 23.71 0.754 0.486 1.061 4.199 26.74 0.715 0.421
Ours 600 1.105 1.908 23.88 0.763 0.492 1.041 4.179 27.00 0.720 0.425
@@
We compare the separation performance of our HSM-TSS with several text-queried audio separation methods, LASSNet [liu2022separate], AudioSep [liu2024separate], CLIPSep [dong2022clipsep], OmniSep [cheng2024omnisep] and BiModalSep [mahmud2024weakly], which are all frequency-domain mask-based approaches. We also compare with a generative model, FlowSep [yuan2025flowsep]. LASS-Net employs a pre-trained BERT as the text query encoder and ResUNet as the separation model, while AudioSep further integrates a CLAP text encoder as the query network and trains on a much larger dataset, yielding substantial performance gains. CLIPSep adopts CLIP as the query encoder and a Sound-of-Pixels (SOP)-based separation model, trained on approximately 500 hours of noisy audio-visual data from the VGGSound [chen2020vggsound] dataset using hybrid vision-text supervision. BiModalSep introduces a weakly-supervised approach and leverages bi-modal semantic similarity via CLAP to align single-source language prompts with audio predictions, achieving robust separation without curated single-source data. OmniSep leverages multi-modal queries flexibly through a omni-modal encoder ImageBind [girdhar2023imagebind]. FlowSep utilizes rectified flow matching for separation. We evaluate their performance using their official open-source models on our benchmarks.
@@
To show the effectiveness of our hierarchical modeling, we also compare with a single-separation-stage variant of our HSM-TSS, termed “Ours_local” in Table II and Table III. It removes the first global-semantic separation stage and only leverages the local-semantic separation stage. The text encoder is FLAN-T5 without Q-Audio in this variant. Additionally, we present an ablation study by evaluating another single-separation-stage variant which leverages Q-Audio text encoder, noted as “Ours_local*”. We provide results for not only “extraction” as used in six compared methods but also “removal” task types that produce the same target audios.
@@
As shown in Table II and Table III, our variant “Ours_local” outperforms baselines in several metrics by a large margin. With the introduction of another global-semantic separation, the proposed HSM-TSS, noted as “Ours”, outperforms both baseline methods and “Ours_local” across various metrics. When compared to other methods, LASSNet, CLIPSep, AudioSep, OmniSep, FlowSep and BiModalSep, our approach consistently achieves higher scores in the three objective metrics, especially for the 3 Sets benchmark, highlighting its effectiveness in separating target audio events from complex mixtures. In contrast to the single-separation-stage setting, our dual-stage separation allows for progressive refinement of audio features from coarse to fine, leading to more precise separation of target sounds, e.g. KL divergence decreases from 0.999 to 0.924 on 3 Sets. We can also see that our model with “removal” instruction achieves comparable scores with “extraction”, showing the bidirectional operation capability of our method.
@@
Table II and Table III also show the evaluation of semantic correctness by CLAP score and AFSim. For “removal” setting, we use the same audio description as “extraction” to calculate the CLAP score. We can see that our method consistently outperforms other methods and the single-separation-stage variants in both AFSim and CLAP scores, showing its superior capability to follow instructions with diverse audio descriptions. Our method has good arbitrary instruction following capability, as shown in our demo page [^3] https://hsm-tss.github.io.
It is also observed that the impact of using the Q-Audio text encoder in the single-separation-stage model is marginal. The performance of “Ours_local*” and “Ours_local” remains nearly identical across multiple evaluation metrics, and both are consistently inferior to “Ours”. This is reasonable as Q-Audio is designed to align text and audio in the shared space, but in local-semantic-separation stage, the audio features are from AudioMAE instead of Q-Audio. Further extracting text feature from FLAN-T5 does not alter its semantic representativeness.
@@ V-B Zero-shot performance on unseen datasets
TABLE IV: Evaluation on zero-shot MUSIC dataset. “Ours_local” denotes the single-separation-stage variant using only local-semantic separation.
Model KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)}
GT - - - - 0.490
LASS-Net 4.180 2.066 13.23 0.503 0.126
CLIPSep 2.179 2.554 14.07 0.657 0.372
BiModalSep 2.607 2.250 14.35 0.671 0.374
AudioSep 0.535 1.624 22.01 0.804 0.342
OmniSep 1.200 2.161 16.91 0.716 0.383
FlowSep 1.659 1.706 18.20 0.733 0.437
@@
Ours_local 0.585 1.400 23.22 0.799 0.463
Ours 0.501 1.375 23.85 0.811 0.470
V-C Ablation study on local-semantic representation
To further investigate the contribution of AudioMAE as the local-semantic representation, we conduct additional ablation experiments under the single-separation-stage setting without a global-semantic separation. As shown in Table V, we compare “Ours_local” with two other variants, “Ours_local_CLAP” and “Ours_acoustic”. In “Ours_local_CLAP”, we replace AudioMAE with the last-layer local features from the CLAP audio encoder [wu2023large] before pooling. Although CLAP features exhibit strong semantic alignment with text, they tend to compromise local acoustic details for reconstruction. The “Ours_acoustic” omits AudioMAE as the local semantic separation and directly utilizes neural codec embeddings for separation by a transformer. While these embeddings retain detailed acoustic information, they lack semantic abstraction, which makes it difficult to separate target signals following text instructions. As shown in Table V, both alternatives yield much inferior results compared to “Ours_local” using AudioMAE across all metrics. These findings confirm that neither approach can capture fine-grained acoustic details and semantic relevance as effectively as AudioMAE.
TABLE V: Ablation results of AudioMAE on 3 Sets. “Ours_local” denotes the single-separation-stage variant using AudioMAE for local-semantic separation. “Ours_local_CLAP” denotes the single-separation-stage variant using CLAP encoder feature for local-semantic separation. “Ours_acoustic” denotes the single-acoustic-separation variant performing separation on acoustic features only without AudioMAE.
Model KL(){($\downarrow$)} LSD(){($\downarrow$)} PSNR(){($\uparrow$)} AFSim(){($\uparrow$)} CLAP(){($\uparrow$)}
GT - - - - 0.481
Extraction
Ours_acoustic (w/o AudioMAE) 2.024 3.106 16.59 0.601 0.293
Ours_local_CLAP 1.901 3.002 17.95 0.634 0.310
Ours_local 0.999 2.878 25.50 0.744 0.420
V-D Impact of source overlap
To assess the robustness of the proposed model under varying levels of difficulty, we perform an ablation study by adjusting the overlap ratio between two source signals. Figure 5 illustrates the separation results in terms of KL divergence and AFSim across different overlap levels.
It can be seen that as the overlap ratio increases, the separation task becomes more challenging. However, our model consistently outperforms the baselines. More importantly, the performance margin becomes more pronounced under higher overlap ratios, where our method achieves the largest relative improvement over both “Ours_local” and “AudioSep”. These findings demonstrate that the proposed framework is particularly effective in handling complex mixtures with heavily overlapped sources, highlighting its robustness in real-world acoustic scenarios.
Figure 5: The trends of KL and AFSim with varying source overlap ratios.
V-E Visualization
V-E1 t-SNE visualization
To show how the global-semantic separation performs, we visualize the extracted features from this stage with t-SNE [van2008visualizing]. In Figure 6, each color shows a sound event class and we present the ground-truth global audio semantic feature and the separated output with different markers. It can be observed that our global-semantic separation model effectively extracts target audio features with good clustering.
@@ V-E Visualization
V-E2 Visualization of attention maps
@@
V-E3 Visualization of separation outputs
@@
V-F Semantic representation evaluation
V-F1 Global semantic feature evaluation
To evaluate the cross-modal alignment and semantic representation capbility of our Q-Audio, we follow the settings of LAION-CLAP [wu2023large] on AudioCaps and Clotho datasets. Table VI shows the text-to-audio retrieval results. Compared to the widely used LAION-CLAP [wu2023large] and MSCLAP [elizalde2023clap], Q-Audio achieves better performance, underscoring its capability in audio-language modeling. Table VII shows the audio captioning evaluation, where a linear mapping layer and a pretrained GPT2 [radford2019language] are used as the downstream model to generate audio captions. Here we only train the downstream model and keep audio feature extraction frozen. Although our Q-Audio is not pretrained on the captioning task as CLAP models, it still achieves higher scores on the three metrics, showing its excellent semantic representation capability.
TABLE VI: Text-to-audio retrieval performance